# neurips_data_2024.bib

### v1

@@ Appendix C Experimental details
For reproducibility, we provide a Python package extending the dynamic graph learning library DyGLib [^3] https://github.com/yule-BUAA/DyGLib (MIT License) [BetterDynamicGraph2023yu] as a supplement, including a bash script to run the experiments. The code will be made publicly available on GitHub after acceptance of the paper.
[…]
@@
We use 9 state-of-the-art dynamic graph learning models and baselines (JODIE [PredictingDynamicEmbedding2019kumar], DyRep [DyRepLearningRepresentations2019trivedi], TGAT [InductiveRepresentationLearning2020xu], TGN [TemporalGraphNetworks2020rossi], CAWN [InductiveRepresentationLearning2021wang], EdgeBank [BetterEvaluationDynamic2022poursafaei], TCL [TCLTransformerbasedDynamic2021wanga], GraphMixer [WeReallyNeed2023cong] and DyGFormer [BetterDynamicGraph2023yu]). The neural-network-based approaches (all except EdgeBank) are trained five times for 100 epochs using the Adam optimizer with a learning rate of $0.0001$. An early-stopping strategy with a patience of 5 is employed to avoid overfitting. For training and validation, a batch size of 200 is used. The training, validation and test sets of each dataset contain 70%, 15% and 15% of the edges, respectively. The sets are split based on time, i.e., the training set contains the edges that occurred first while the test set comprises the most recent edges.
@@
(a) Hyperparameters for neighborhood sampling-based models. $n-\text{Neighbors}$ is the number of sampled neighbors using the specified neighbor sampling strategy. $n-\text{Layers}$ is the number of transformer layers (TCL), the number of MLP-Mixer layers (for GraphMixer) or the number of GNN layers otherwise.
@@
Appendix D Detailed AUC-ROC and average precision results
@@
Here, we provide detailed tabulated results for all models’ AUC-ROC and average precision performance across five runs, including standard deviations.
AUC-ROC
Table 7: Average AUC-ROC performance over five runs for the test set of the continuous-time datasets from [BetterEvaluationDynamic2022poursafaei, BetterDynamicGraph2023yu], including standard deviations.
@@
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
@@
LastFM 78.0 ± 0.7 71.7 ± 1.1 80.7 ± 2.4 68.4 ± 0.7 68.1 ± 0.3 78.2 ± 0.0 64.3 ± 6.1 65.9 ± 1.7 78.9 ± 0.6
Myket 64.0 ± 2.1 64.2 ± 2.7 61.1 ± 2.6 57.6 ± 0.4 32.5 ± 0.4 51.9 ± 0.0 58.4 ± 2.0 59.5 ± 0.4 32.8 ± 1.0
Social 91.4 ± 2.1 92.7 ± 0.5 91.7 ± 3.3 92.6 ± 0.5 87.7 ± 0.1 85.8 ± 0.0 95.2 ± 0.2 94.1 ± 0.2 97.3 ± 0.1
@@
Table 8: Average AUC-ROC performance over five runs for the test set of the discrete-time datasets from [BetterEvaluationDynamic2022poursafaei, BetterDynamicGraph2023yu], including standard deviation.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
@@
Flights 69.5 ± 2.2 69.0 ± 1.0 68.8 ± 1.6 72.6 ± 0.2 65.0 ± 1.4 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.9 ± 1.1
Cont. 95.5 ± 0.6 95.4 ± 0.2 96.1 ± 0.8 95.4 ± 0.3 83.3 ± 0.0 92.2 ± 0.0 94.1 ± 0.8 94.1 ± 0.2 97.1 ± 0.0
Average precision
@@
MOOC 82.4( $\downarrow$1.1%) 76.9( $\downarrow$0.3%) 86.3( $\downarrow$0.8%) 82.7( $\downarrow$2.1%) 72.3( $\downarrow$1.7%) 59.1( $\downarrow$2.7%) 74.9( $\downarrow$4.6%) 74.4( $\downarrow$4.5%) 82.1( $\downarrow$0.4%) 2.0%±1.6%
Wiki. 84.1( $\downarrow$0.1%) 80.9( $\uparrow$0.1%) 88.5( $\downarrow$0.4%) 87.5( $\downarrow$0.4%) 75.1( $\uparrow$0.1%) 73.3( $\uparrow$0.2%) 89.2( $\downarrow$0.4%) 90.8( $\downarrow$0.4%) 83.1( $\uparrow$0.0%) 0.2%±0.2%
@@
Myket 64.5( $\uparrow$1.6%) 63.1( $\uparrow$0.9%) 62.8( $\uparrow$1.2%) 57.9( $\uparrow$1.4%) 46.6( $\uparrow$3.3%) 51.9( $\uparrow$1.3%) 58.9( $\uparrow$1.1%) 60.0( $\uparrow$1.5%) 46.1( $\uparrow$3.3%) 1.7%±0.9%
Social 89.4( $\uparrow$1.2%) 91.9( $\uparrow$0.3%) 93.9( $\uparrow$0.8%) 95.0( $\uparrow$0.2%) 85.6( $\downarrow$0.6%) 79.7( $\downarrow$1.1%) 96.1( $\downarrow$0.1%) 95.8( $\uparrow$0.4%) 97.7( $\uparrow$0.3%) 0.6%±0.4%
@@
UN Tr. 52.7( $\downarrow$10.6%) 49.4( $\downarrow$16.6%) 53.2( $\downarrow$9.7%) 59.1( $\uparrow$2.0%) 59.2( $\uparrow$2.4%) 79.0( $\downarrow$2.6%) 57.5( $\uparrow$2.3%) 65.8( $\uparrow$3.3%) 67.1( $\uparrow$4.4%) 6.0%±5.1%
Can. P. 52.1( $\downarrow$1.3%) 61.0( $\downarrow$1.3%) 69.9( $\uparrow$2.0%) 70.8( $\uparrow$4.5%) 68.3( $\uparrow$6.9%) 59.4( $\downarrow$6.8%) 68.2( $\uparrow$6.2%) 80.9( $\uparrow$4.9%) 83.2( $\downarrow$14.3%) 5.4%±4.0%
Flights 65.2( $\downarrow$2.2%) 63.9( $\downarrow$4.3%) 68.3( $\downarrow$0.0%) 73.5( $\uparrow$1.1%) 64.7( $\uparrow$1.3%) 70.4( $\downarrow$0.2%) 71.0( $\uparrow$0.4%) 71.9( $\uparrow$1.0%) 68.9( $\downarrow$0.1%) 1.2%±1.4%
Cont. 94.0( $\downarrow$0.2%) 95.8( $\uparrow$0.4%) 97.0( $\uparrow$0.8%) 96.8( $\uparrow$0.8%) 88.2( $\uparrow$4.4%) 89.4( $\uparrow$0.6%) 96.6( $\uparrow$2.1%) 95.7( $\uparrow$1.6%) 98.3( $\uparrow$0.6%) 1.3%±1.3%
$\mu\pm\sigma$ 6.9%±8.5% 10.6%±10.4% 10.8%±12.5% 1.6%±1.5% 3.1%±2.4% 1.8%±2.6% 2.1%±2.2% 1.9%±1.8% 4.0%±5.4%
Table 11: Mean average precision performance over five runs for the test set of the continuous-time datasets from [BetterEvaluationDynamic2022poursafaei, BetterDynamicGraph2023yu], including standard deviations.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
@@
Table 12: Mean average precision performance over five runs for the test set of the discrete-time datasets from [BetterEvaluationDynamic2022poursafaei, BetterDynamicGraph2023yu], including standard deviations.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer

### v2

@@ Appendix F Experimental Details
For reproducibility, we provide a GitHub repository [^3] https://github.com/M-Lampert/DyGLib extending the dynamic graph learning library DyGLib [^4] https://github.com/yule-BUAA/DyGLib (MIT License) (Yu et al., 2023), including a bash script to run the experiments.
We use the best hyperparameters reported by Yu et al. (2023) and, for completeness, list these hyperparameters for the 13 datasets used by Yu et al. (2023) below. However, the Myket dataset (Loghmani & Fazli, 2023) was not included in the study. Therefore, for Myket, we use each method’s default parameters as suggested by the respective authors.
@@
We use 9 state-of-the-art dynamic graph learning models and baselines (JODIE (Kumar et al., 2019), DyRep (Trivedi et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), CAWN (Wang et al., 2021b), EdgeBank (Poursafaei et al., 2022), TCL (Wang et al., 2021a), GraphMixer (Cong et al., 2023) and DyGFormer (Yu et al., 2023)). The neural-network-based approaches (all except EdgeBank) are trained five times for 100 epochs using the Adam optimizer with a learning rate of $0.0001$. An early-stopping strategy with a patience of 5 is employed to avoid overfitting. For training and validation, a batch size of 200 is used. The training, validation and test sets of each dataset contain 70%, 15% and 15% of the edges, respectively. The sets are split based on time, i.e., the training set contains the edges that occurred first while the test set comprises the most recent edges.
@@
AMD Epyc 7543 (3rd Gen)
AMD Epyc 7763 (3rd Gen)
AMD Epyc 7713 (3rd Gen)
Intel Xeon Platinum 8480+ (4th Gen)
@@
(a) Hyperparameters for neighborhood sampling-based models. $n_{\text{Neighbors}}$ is the number of sampled neighbors using the specified neighbor sampling strategy. $n_{\text{Layers}}$ is the number of transformer layers (for TCL), the number of MLP-Mixer layers (for GraphMixer) or the number of GNN layers otherwise.
@@
Appendix G Detailed AUC-ROC and Average Precision Results
@@
Here, we provide the evaluation runtime of all models for dynamic link prediction and forecasting averaged across five runs and detailed tabulated results for all models’ AUC-ROC and average precision performance across five runs, including standard deviations. We further include visualisations investigating the causes of the large performance differences between dynamic link prediction and forecasting.
G.1 Runtime
Table 11 shows the runtime during the evaluation of all models for dynamic link prediction and forecasting. The runtime of both approaches is generally comparable with both approaches being faster than the other on some datasets. For continuous-time datasets, we can see that the runtime is often comparable and large differences are caused by specific characteristics of the datasets. The model evaluation on UCI, for example, is faster using the batch-based approach because it consistently evaluates batches of size $b=200$ while the time windows are mostly smaller which leads to less parallelization. This is because the duration of each window was chosen so that the average window size across the whole dataset is comparable to the chosen batch size of $b=200$. For this specific dataset, there are large bursts of activity in the first third of the total dataset duration and low activity afterwards. This leads to small window sizes for the temporal edges of the test set (see Figure 5). For discrete-time datasets, each snapshot contains more than 200 edges for most datasets leading to better parallelization for link forecasting and, thus, smaller runtime.
Table 11: Mean and standard deviation in seconds of the evaluation runtime of all models for dynamic link prediction and forecasting averaged across five runs.
@@ G.1 Runtime
Dataset Approach JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Enron Forec. 3.42 ± 0.26 4.15 ± 0.30 4.20 ± 0.31 36.16 ± 0.35 67.98 ± 0.76 9.00 ± 0.11 7.76 ± 0.26 12.44 ± 0.33 43.58 ± 0.43
Pred. 3.47 ± 1.57 4.16 ± 1.99 4.29 ± 1.92 37.26 ± 2.03 73.06 ± 1.38 7.74 ± 0.09 7.61 ± 1.18 12.65 ± 1.24 47.79 ± 1.69
UCI Forec. 6.84 ± 0.47 8.40 ± 0.50 8.44 ± 0.50 37.72 ± 0.44 97.97 ± 1.58 4.56 ± 0.17 11.36 ± 0.30 13.90 ± 0.31 23.85 ± 0.39
Pred. 2.66 ± 1.02 3.08 ± 0.72 3.09 ± 0.72 24.58 ± 0.83 88.00 ± 2.72 1.15 ± 0.02 4.86 ± 0.97 8.03 ± 0.99 18.29 ± 0.78
MOOC Forec. 55.41 ± 0.52 59.23 ± 0.37 58.56 ± 0.38 210.16 ± 0.28 539.03 ± 2.59 264.74 ± 2.79 89.27 ± 0.20 132.41 ± 0.42 226.80 ± 1.38
Pred. 61.10 ± 0.56 64.53 ± 0.43 65.49 ± 0.35 230.79 ± 1.74 555.57 ± 3.37 301.55 ± 5.18 95.60 ± 1.19 140.46 ± 1.36 254.47 ± 0.93
Wiki. Forec. 15.16 ± 0.73 16.29 ± 0.79 16.46 ± 0.65 44.58 ± 0.88 90.17 ± 1.10 10.27 ± 0.46 17.19 ± 0.70 19.80 ± 0.71 50.72 ± 1.13
Pred. 13.67 ± 1.08 14.63 ± 0.84 13.91 ± 0.91 45.38 ± 1.74 89.25 ± 1.72 9.46 ± 0.43 15.99 ± 1.44 19.01 ± 1.46 49.27 ± 1.50
LastFM Forec. 271.34 ± 0.97 281.58 ± 1.81 338.74 ± 1.73 985.54 ± 34.93 3316.72 ± 21.18 $1$ 18.08 578.48 ± 3.48 702.40 ± 1.33 1037.08 ± 8.49
Pred. 319.89 ± 1.76 340.54 ± 12.63 400.65 ± 11.46 1180.93 ± 83.27 3472.23 ± 79.24 1723.99 ± 28.39 674.63 ± 12.96 832.52 ± 3.00 1193.28 ± 3.40
Myket Forec. 467.24 ± 3.94 465.55 ± 6.40 445.55 ± 11.19 680.26 ± 15.71 851.62 ± 21.51 479.92 ± 6.43 507.72 ± 7.10 602.63 ± 7.31 643.33 ± 8.06
Pred. 425.36 ± 6.72 424.50 ± 7.38 403.12 ± 9.27 701.21 ± 8.81 876.41 ± 17.09 445.81 ± 5.62 475.26 ± 8.32 566.69 ± 19.39 628.24 ± 7.03
Social Forec. 408.85 ± 1.89 424.96 ± 3.62 419.62 ± 2.44 1208.04 ± 12.30 4020.21 ± 4.69 1090.46 ± 1.08 709.80 ± 2.45 899.58 ± 2.20 1025.54 ± 2.26
Pred. 500.15 ± 3.65 509.14 ± 2.35 510.64 ± 3.11 1544.19 ± 2.59 4308.14 ± 8.00 1379.71 ± 0.74 955.07 ± 1.34 1203.01 ± 1.67 1263.80 ± 4.42
Reddit Forec. 120.79 ± 0.77 124.65 ± 0.74 131.35 ± 2.35 543.41 ± 1.25 561.97 ± 2.74 174.94 ± 0.50 161.26 ± 1.12 213.45 ± 0.73 248.81 ± 1.26
Pred. 120.22 ± 0.46 124.28 ± 0.41 129.58 ± 0.67 585.98 ± 53.62 575.36 ± 2.94 191.38 ± 0.59 167.83 ± 1.06 227.86 ± 2.07 254.34 ± 1.45
UN V. Forec. 4.25 ± 0.30 9.37 ± 0.28 20.11 ± 0.36 161.34 ± 1.82 5804.24 ± 9.62 6.88 ± 0.05 40.01 ± 0.29 831.89 ± 9.65 137.97 ± 0.97
Pred. 158.35 ± 0.46 161.49 ± 0.70 174.23 ± 0.68 560.17 ± 4.39 5599.27 ± 92.68 684.81 ± 8.57 310.53 ± 0.78 1223.42 ± 8.09 406.54 ± 0.94
US L. Forec. 1.01 ± 0.26 1.33 ± 0.31 1.32 ± 0.31 13.25 ± 0.30 32.43 ± 0.55 0.05 ± 0.01 3.19 ± 0.25 3.63 ± 0.26 18.67 ± 0.26
Pred. 2.55 ± 1.82 2.93 ± 1.76 2.84 ± 1.77 19.40 ± 0.82 42.72 ± 1.06 0.71 ± 0.02 5.73 ± 0.98 6.87 ± 1.25 26.85 ± 1.21
UN Tr. Forec. 1.95 ± 0.26 4.13 ± 0.24 3.97 ± 0.12 271.15 ± 3.19 2813.05 ± 2.68 1.83 ± 0.06 19.42 ± 0.26 507.83 ± 14.74 84.71 ± 0.52
Pred. 34.63 ± 0.42 37.36 ± 0.44 36.37 ± 0.44 466.44 ± 2.36 2992.96 ± 384.16 161.85 ± 0.63 90.81 ± 0.43 719.00 ± 70.07 213.38 ± 1.02
Can. P. Forec. 0.56 ± 0.22 2.00 ± 0.24 1.99 ± 0.24 60.15 ± 0.31 1034.22 ± 1.06 0.11 ± 0.02 4.11 ± 0.21 20.65 ± 0.30 464.45 ± 25.61
Pred. 2.38 ± 0.46 3.68 ± 0.54 3.80 ± 0.56 90.50 ± 1.56 1242.03 ± 185.41 2.02 ± 0.03 7.64 ± 0.63 30.36 ± 1.72 380.33 ± 2.02
Flights Forec. 343.75 ± 10.64 346.74 ± 9.28 341.56 ± 9.66 357.77 ± 5.58 10363.33 ± 2098.15 21.36 ± 0.44 68.14 ± 0.78 1611.70 ± 33.57 994.11 ± 7.16
Pred. 1567.52 ± 79.75 1550.22 ± 94.86 1581.06 ± 47.29 2359.88 ± 54.11 3420.00 ± 57.27 2132.99 ± 103.26 1933.46 ± 84.54 2113.83 ± 20.74 2701.10 ± 31.46
Cont. Forec. 656.31 ± 2.85 673.36 ± 3.62 683.82 ± 2.96 1966.25 ± 10.58 3878.32 ± 11.95 2646.55 ± 37.98 1139.57 ± 4.54 1378.58 ± 4.22 1518.44 ± 12.00
Pred. 938.48 ± 6.81 940.55 ± 7.94 953.46 ± 8.02 2528.31 ± 34.70 4358.78 ± 21.10 3854.40 ± 32.23 1523.29 ± 26.55 1848.28 ± 40.53 1873.09 ± 11.28
G.2 AUC-ROC
The following reports the complete results (including standard deviations) of the experiments conducted in Section 4.
Table 12: Average AUC-ROC performance over five runs for the test set of the continuous-time datasets, including standard deviations. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
@@ G.2 AUC-ROC
LastFM 78.0 ± 0.7 71.7 ± 1.1 80.7 ± 2.4 68.4 ± 0.7 68.1 ± 0.3 78.2 ± 0.0 64.3 ± 6.0 65.9 ± 1.7 78.9 ± 0.6
Myket 64.0 ± 2.1 64.2 ± 2.7 61.1 ± 2.6 57.6 ± 0.4 32.5 ± 0.4 52.0 ± 0.0 58.4 ± 2.0 59.5 ± 0.4 32.8 ± 1.0
Social 91.3 ± 2.1 92.7 ± 0.5 91.7 ± 3.3 92.6 ± 0.5 87.7 ± 0.1 85.8 ± 0.0 95.2 ± 0.2 94.1 ± 0.2 97.3 ± 0.1
@@
Avg. Rank 3.1 5.1 3.4 5.8 7.6 6.1 5.9 4.1 3.9
Table 13: Average AUC-ROC performance over five runs for the test set of the discrete-time datasets, including standard deviation. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
@@
Flights 69.5 ± 2.2 69.0 ± 1.0 68.8 ± 1.6 72.6 ± 0.2 66.0 ± 1.7 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.8 ± 1.0
Cont. 95.5 ± 0.5 95.4 ± 0.2 96.1 ± 0.8 95.4 ± 0.3 83.3 ± 0.0 92.2 ± 0.0 94.5 ± 0.5 94.1 ± 0.2 97.1 ± 0.0
Avg. Rank 5.2 5.3 4.3 5.3 7.0 4.7 6.3 3.7 3.2
G.3 Average precision
Average precision scores as an additional metric to evaluate experiments in Section 4. The results are comparable to the results using the AUC-ROC score.
@@ G.3 Average precision
MOOC 82.4( $\downarrow$1.1%) 76.9( $\downarrow$0.3%) 86.3( $\downarrow$0.8%) 82.7( $\downarrow$2.1%) 72.3( $\downarrow$1.7%) 59.0( $\downarrow$2.7%) 74.9( $\downarrow$4.6%) 74.4( $\downarrow$4.5%) 82.1( $\downarrow$0.4%) 2.0%±1.6%
Wiki. 84.1( $\downarrow$0.1%) 80.9( $\uparrow$0.1%) 88.5( $\downarrow$0.4%) 87.5( $\downarrow$0.4%) 75.1( $\uparrow$0.1%) 73.3( $\uparrow$0.3%) 89.2( $\downarrow$0.4%) 90.8( $\downarrow$0.4%) 83.1( $\uparrow$0.0%) 0.2%±0.2%
@@
Myket 64.5( $\uparrow$1.6%) 63.1( $\uparrow$0.9%) 62.8( $\uparrow$1.2%) 57.9( $\uparrow$1.4%) 46.6( $\uparrow$3.3%) 51.9( $\uparrow$1.3%) 58.9( $\uparrow$1.0%) 60.0( $\uparrow$1.5%) 46.1( $\uparrow$3.3%) 1.7%±0.9%
Social 89.4( $\uparrow$1.3%) 91.9( $\uparrow$0.3%) 93.9( $\uparrow$0.8%) 95.0( $\uparrow$0.2%) 85.6( $\downarrow$0.6%) 79.7( $\downarrow$1.1%) 96.1( $\downarrow$0.1%) 95.8( $\uparrow$0.4%) 97.7( $\uparrow$0.3%) 0.6%±0.4%
@@
UN Tr. 52.7( $\downarrow$10.6%) 49.4( $\downarrow$16.6%) 53.2( $\downarrow$9.7%) 58.9( $\uparrow$1.7%) 59.2( $\uparrow$2.4%) 79.0( $\downarrow$2.6%) 57.5( $\uparrow$2.3%) 65.8( $\uparrow$3.3%) 67.1( $\uparrow$4.4%) 6.0%±5.2%
Can. P. 52.1( $\downarrow$1.3%) 61.0( $\downarrow$1.3%) 69.9( $\uparrow$2.0%) 70.8( $\uparrow$4.5%) 68.3( $\uparrow$6.9%) 59.4( $\downarrow$6.8%) 68.2( $\uparrow$6.2%) 80.9( $\uparrow$4.8%) 83.2( $\downarrow$14.3%) 5.3%±4.0%
Flights 65.2( $\downarrow$2.2%) 63.9( $\downarrow$4.3%) 68.3( $\downarrow$0.0%) 73.5( $\uparrow$1.1%) 65.5( $\uparrow$0.8%) 70.4( $\downarrow$0.2%) 71.0( $\uparrow$0.4%) 71.9( $\uparrow$1.0%) 69.1( $\uparrow$0.7%) 1.2%±1.3%
Cont. 94.0( $\downarrow$0.2%) 95.8( $\uparrow$0.3%) 97.0( $\uparrow$0.8%) 96.8( $\uparrow$0.8%) 88.2( $\uparrow$4.4%) 89.4( $\uparrow$0.6%) 96.6( $\uparrow$1.8%) 95.7( $\uparrow$1.6%) 98.3( $\uparrow$0.6%) 1.2%±1.3%
$\mu\pm\sigma$ 6.9%±8.5% 10.6%±10.4% 10.8%±12.5% 1.6%±1.5% 3.0%±2.5% 1.8%±2.6% 2.1%±2.2% 1.9%±1.8% 4.1%±5.3%
Table 16: Mean average precision performance over five runs for the test set of the continuous-time datasets, including standard deviations. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
@@
Avg. Rank 3.6 6.2 3.6 5.1 7.1 7.4 4.9 3.5 3.5
Table 17: Mean average precision performance over five runs for the test set of the discrete-time datasets, including standard deviations. For discrete- and continuous-time datasets, the average rank of each model is reported in the last row, respectively.
Eval. Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer